# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ VII-B3 Distributed backdoor attack
Backdoor attacks performed by several different collaborative attackers, i.e. distributed backdoor attacks, can reach very successful rates and are very difficult to be detected. Based on a dynamic norm clipping approach, Guo et al. [154] proposed a security system to defend against distributed backdoor attacks in federated learning. The proposed system is evaluated with the following four public datasets: MNIST, FMNIST, CIFAR-10, and Tiny Imagenet. The results show that the proposed system can reduce the attack access rate by 84.23% compared to the state of the art. With the distributed backdoor attack, there are attackers who constitute a group with identical malicious purposes. More precisely, all attackers conduct the backdoor attack on their local models independently using their specific local backdoor trigger ${Trig}_{i}$ and the same target label ${Label}_{target}$. The attacker’s corresponding goal $i$ is defined as follows :
@@
Where $w$ represents the global model parameters, ${Data}^{i}_{infected}$ is the infected image dataset of attacker $i$, ${Data}^{i}_{uninfected}$ is the uninfected image dataset of attacker $i$, $x^{i}_{j}$ represents input from the local dataset, and $y^{i}_{j}$ represent the corresponding label sampled from the local dataset. The objective of each attacker is to maximize the misclassification rate of the global model on a specific target label ${Label}_{target}$ when presented with inputs that contain a specific backdoor trigger ${Trig}_{i}$. The goal $G_{i}\left(w\right)$ of attacker $i$ is defined as the sum of two terms: the first term represents the misclassification rate of the infected images from attacker $i$ on the target label ${Label}_{target}$, while the second term represents the misclassification rate of the uninfected images from attacker $i$ on their true labels. The function $f(w;x,y)$ represents the output of the global model with parameters $w$ on input $x$ and label $y$. The double semicolon notation $w;;x^{i}_{j}+{Trig}_{i}$ indicates that the backdoor trigger ${Trig}_{i}$ is added to the input $x^{i}_{j}$ before it is fed into the global model.
@@ VII-B4 Clean label backdoor attack
Ning et al. [178] proposed a black box clean label backdoor attack, which requires zero knowledge of the target model. In order to detect and eliminate, the authors propose the following two methods: 1) Supervised poison sample detection and 2) Unsupervised poison sample detection. For Backdoor injection, the adversary summarized another dataset by combining the clean database with a small part of the poisoned dataset as follows:
@@
The adversary uses this dataset during the training process and verifies that other properties are not influenced by a combined dataset. The backdoor injection attack involves adding malicious data to the training dataset, which can be used to exploit the model during inference. In this case, the adversary has combined a clean dataset with a small part of the poisoned dataset to create a new dataset for training. The poisoned dataset contains inputs and labels that have been specifically chosen to trigger a backdoor in the model. To train the model using this dataset, the adversary minimizes a loss function that includes both clean and poisoned data. The goal is to find the parameters of the model that minimize the total loss on both datasets. The loss function is typically based on cross-entropy, which measures the difference between the predicted probabilities and the true labels.
@@
$model$ represents the parameters of the model. $Loss(model,\left(x^{i},y^{j}\right))$ represents the loss incurred by the model when making a prediction for input $x^{i}$ with label $y^{j}$ from the poisoned dataset. $Loss(model,\left(x^{i^{\prime}},y^{j^{\prime}}\right))$ represents the loss incurred by the model when making a prediction for input $x^{i^{\prime}}$ with label $y^{j^{\prime}}$ from the clean dataset. $n$ is the size of the poisoned dataset. $m$ is the size of the clean dataset.
@@ VII-B5 Pixel-space backdoor attack
Pixel-space backdoor attacks are a type of adversarial attack where an attacker modifies the pixel values of an image in a way that is imperceptible to the human eye but causes misclassification by the target model. Arshad et al.’s pixel-space backdoor attack [209] is a type of data poisoning attack, where the attacker injects malicious samples into the training dataset to cause the model to learn a backdoor trigger. In this attack, the attacker flips certain bits of the pixel values of the training images, which can be done without significantly changing the appearance of the image. These modified images are then inserted into the training dataset, along with their original labels. During training, the model learns to associate the backdoor trigger with the target label, allowing the attacker to trigger the attack by adding the trigger to any input image. Arshad et al. evaluated their backdoor attack strategy on the MNIST digit dataset, which consists of 28x28 Grey-scale images of handwritten digits. They found that their attack was successful in 90% of the cases, meaning that the target model misclassified the poisoned images with the intended target label. Therefore, it’s important to note that the success of a backdoor attack depends on several factors, including the size and quality of the training dataset, the choice of the backdoor trigger, and the strength of the attacker’s evasion techniques. Therefore, while Arshad et al.’s attack achieved a high success rate on the MNIST dataset, it may not be as effective on other datasets or against more robust models with stronger defenses against adversarial attacks.
TABLE XIV: Threat models of adversarial examples generation.
Attacks examples Attacks for the specific tasks White/black/Grey box Adversarial Example Generation Methodes
@@
Trained in a white-box or black-box manner
@@
Black/Grey-box attack
@@
Black/Grey-box attack
@@
Black/Grey-box attack
@@
Black/Grey/White-box attack
@@
Black/Grey/White-box attack
@@
Black/Grey/White-box attack
@@ VII-C Adversarial examples
Adversarial examples refer to inputs specifically crafted to deceive machine learning models. Different types of adversarial attacks are presented, including black-box attacks, white-box attacks, and Grey-box attacks. Akhtar and Mian [232] proposed a classification of adversarial examples on deep learning into two categories: 1) Attacks for Classification and 2) Attacks Beyond Classification/Recognition. The Attacks for Classification include miscellaneous attacks, adversarial transformation networks (ATNs), Houdini, Upset and ANGRI, Universal Adversarial Perturbations, Deepfool, Carlini and Wagner Attacks (C&W), One Pixel Attack, Jacobian-Based Saliency Map Attack (JSMA), Basic & Least-Likely-Class Iterative Methods, Fast Gradient Sign Method (FGSM), and Box-Constrained L-BFGS. Attacks Beyond Classification/Recognition include Attacks on Face Attributes, Attacks on Semantic Segmentation & Object Detection, Attacks on Deep Reinforcement Learning, Attacks on Recurrent Neural Networks, and Attacks on Autoencoders and Generative Models. Table XIV presents the threat models of adversarial examples generation. Figure 9 presents the adversarial examples in edge learning. Therefore, in our work, we propose a classification of adversarial examples into three categories: 1) White-box adversarial attacks, 2) Black-box adversarial attacks, and 3) Grey-Box adversarial attacks.
@@ VII-C1 White-box adversarial attacks
White-box attacks require full and complete knowledge of the model being targeted, including its training method, architecture, parameter values, and, in most cases, its data training. To manipulate discrete text structure at its one-hot representation, Ebrahimi et al. [169] propose white-box adversarial examples, named HotFlip, which is applied for text classification. The HotFlip system uses the gradient with respect to a one-hot input representation and can produce conflicting patterns with character substitutions ("flips"). The HotFlip also provides access to add and remove transactions by displaying them as character substitution sequences. The performance evaluation with AG’s news dataset and CharCNN-LSTM architecture shows that the adversary selects the flip transaction about 80% of the time and prioritizes suppression over insertion by two-to-one. For fooling deep neural networks, Wang et al. [170] propose white-box attacks based on the integrated gradient. The proposed attacks are evaluated with the following three datasets: CIFAR-10, MNIST, and NSL-KDD. One of the most popular algorithms for generating white-box adversarial examples is the Fast Gradient Sign Method (FGSM) introduced by Goodfellow et al. in 2014 [226]. The FGSM generates adversarial examples by taking the gradient of the loss function with respect to the input and adding a perturbation to the input in the direction of the gradient. The formula for generating adversarial examples using FGSM is as follows:
${\color[rgb]{0,0,0}x_{adv}=x+\epsilon\cdot sign(\nabla_{x}J(\theta,x,y))}$ (6)
@@
VII-C2 Black-box adversarial attacks
Black-box attacks provide a trained model with conflicting examples (during testing) which are produced with no knowledge of the trained model. In particular cases, it is supposed that the adversary has restricted information about the model but certainly does not have knowledge of the model’s parameters. Yang et al. [159] propose a black-box method called TSadv, which attacks deep neural networks on time series. For solving the constrained optimization problem, the TSadv method uses a gradient-free attack strategy. The performance evaluation with the University of California Riverside (UCR) time series dataset shows good results in terms of success rate, the average number of iterations, and mean squared error. Therefore, Biolková and Nguyen [163] propose a black-box adversarial attack in speech recognition. The proposed attack uses a neural predictor that approximates the length of the decision boundary which results in a wrong audio signal transcription. The experiments on the LibriSpeech dataset show that the proposed attack can achieve a better success rate compared to BayesOpt [233] and SignOpt [234]. One of the most popular algorithms for generating black-box adversarial examples is the Boundary Attack introduced by Brendel and Bethge in 2019 [235]. The Boundary Attack generates adversarial examples by iteratively exploring the decision boundary of the model. The formula for generating adversarial examples using Boundary Attack is as follows:
${\color[rgb]{0,0,0}x_{adv}=x+\alpha\cdot\frac{g}{\left|g\right|_{p}}}$ (7)
Where $x$ is the input, $\alpha$ is a small constant that controls the step size, $g$ is the gradient of the loss function with respect to the input, and $\left|g\right|_{p}$ is the $l_{p}$ norm of the gradient.
@@ VII-C3 Grey-box adversarial attacks
Grey-box adversarial attacks involve generating adversarial examples with limited knowledge of the targeted model. The attacker has access to some, but not all, of the model’s internal information. For example, the attacker may know the architecture of the model, but not the weights or biases. The goal of a grey-box attack is to find an adversarial example that can deceive the targeted model into producing incorrect output, even though the attacker does not have full knowledge of the model’s internal workings. This is often done by leveraging the gradients of the model with respect to the input, which can be computed even if the attacker does not know the exact values of the model’s parameters [236]. Grey-box attacks are particularly relevant in real-world scenarios, where attackers may not have access to the full details of a deployed model, but may still be able to mount effective attacks. Defenses against grey-box attacks typically involve techniques such as input sanitization [237] or model hardening, which aim to make it harder for attackers to find effective adversarial examples. One of the most popular algorithms for generating Grey-box adversarial examples is the Transferability Attack introduced by Papernot et al. in 2016 [238]. The Transferability Attack generates adversarial examples by exploiting the transferability property of adversarial examples, which means that adversarial examples generated for one model can also fool another model. The formula for generating adversarial examples using Transferability Attack is as follows:
${\color[rgb]{0,0,0}x_{adv}=x+\epsilon\cdot sign(\nabla_{x}J(\theta_{target},x,y))}$ (8)
This equation represents the process of generating an adversarial example for a given input image $x$, with respect to a target class $\theta_{target}$ and a true label $y$, by adding a small perturbation $\epsilon$ to the original image $x$ in the direction of the sign of the gradient of the target class probability $J(\theta_{target},x,y)$ with respect to the input image $x$.
In most cases, black-box attacks are viewed as more difficult since the attacker has less information to operate with. Nevertheless, Grey box attacks can still be successful if the offensive party has sufficient knowledge of the pattern to produce successful adversarial examples.
@@ VII-D1 Synergetic attack
A synergetic attack consists of a combination of backdoor and adversarial examples against neural network classifiers. Liu et al. [158] proposed a synergetic attack, named AdvTrojan. The AdvTrojan is enabled based only on the model being infected through a backdoor during training, and the entries are thoughtfully disrupted. The AdvTrojan uses two stages that operate in parallel to progressively relocate the targeted input through the decision-making boundary to the adversary’s objective class. In the initial stage, the model is injected with a Trojan backdoor at the time of training. During inference, the Trojan backdoor is enabled by incrementing the relevant target inputs using the pre-specified Trojan trigger. In the second stage, the targeted entry is completed by a prudent mixture of some adversarial disruption and the Trojan horse trigger. In the second stage, the targeted entry is completed by a prudent mixture of the Trojan horse trigger and some adversarial disruption. The adversary disruption enhances the Trojan horse trigger to exchange the entry label in the class targeted by the adversary. In practical terms, the Trojan horse trigger forwards the entry to an artificial position in the space of the entry closest to the boundary of the model’s decision. Afterward, the adversarial disruption makes the terminal push by shifting the forwarded example through the boundary of the decision, initiating the pre-established backdoor. The combined attacks can be defined as follows:

### v2

@@ VII-B3 Distributed backdoor attack
Backdoor attacks performed by several collaborative attackers, i.e., distributed backdoor attacks, can reach very successful rates and are very difficult to detect. A distributed backdoor attack in 6G networks could exploit the advanced and decentralized nature of the system, subtly introducing malicious functionalities without detection. By compromising multiple nodes or devices, the attack can gain control over data flow, causing disruptions or unauthorized data access. Based on a dynamic norm clipping approach, Guo et al. [175] proposed a security system to defend against distributed backdoor attacks in federated learning. The proposed system is evaluated with the following four public datasets: MNIST, FMNIST, CIFAR-10, and Tiny Imagenet. The results show that the proposed system can reduce the attack access rate by 84.23% compared to the state of the art. With the distributed backdoor attack, some attackers constitute a group with identical malicious purposes. More precisely, all attackers conduct the backdoor attack on their local models independently using their specific local backdoor trigger ${Trig}_{i}$ and the same target label ${Label}_{target}$. The attacker’s corresponding goal $i$ is defined as follows :
@@
Where $w$ represents the global model parameters, ${Data}^{i}_{infected}$ is the infected image dataset of attacker $i$, ${Data}^{i}_{uninfected}$ is the uninfected image dataset of attacker $i$, $x^{i}_{j}$ represents input from the local dataset and $y^{i}_{j}$ represent the corresponding label sampled from the local dataset. The objective of each attacker is to maximize the misclassification rate of the global model on a specific target label ${Label}_{target}$ when presented with inputs that contain a specific backdoor trigger ${Trig}_{i}$. The goal $G_{i}\left(w\right)$ of attacker $i$ is defined as the sum of two terms: the first term represents the misclassification rate of the infected images from attacker $i$ on the target label ${Label}_{target}$, while the second term represents the misclassification rate of the uninfected images from attacker $i$ on their true labels. The function $f(w;x,y)$ represents the output of the global model with parameters $w$ on input $x$ and label $y$. The double semicolon notation $w;;x^{i}_{j}+{Trig}_{i}$ indicates that the backdoor trigger ${Trig}_{i}$ is added to the input $x^{i}_{j}$ before it is fed into the global model.
@@ VII-B4 Clean label backdoor attack
A clean-label backdoor attack on 6G could subtly embed malicious triggers in legitimate data during the system’s training phase. Once deployed, adversaries can exploit these triggers to manipulate the 6G network’s behavior, undermining its security and reliability. Such attacks could severely compromise the advanced functionalities and services anticipated in 6G ecosystems. Ning et al. [198] proposed a blue box clean label backdoor attack, which requires zero knowledge of the target model. To detect and eliminate, the authors propose the following two methods: 1) Supervised poison sample detection and 2) Unsupervised poison sample detection. For Backdoor injection, the adversary summarized another dataset by combining the clean database with a small part of the poisoned dataset as follows:
@@
The adversary uses this dataset during training and verifies that a combined dataset does not influence other properties. The backdoor injection attack involves adding malicious data to the training dataset, which can be used to exploit the model during inference. In this case, the adversary has combined a clean dataset with a small part of the poisoned dataset to create a new dataset for training. The poisoned dataset contains inputs and labels specifically chosen to trigger a backdoor in the model. To train the model using this dataset, the adversary minimizes a loss function that includes clean and poisoned data. The goal is to find the parameters of the model that minimize the total loss on both datasets. The loss function is typically based on cross-entropy, which measures the difference between the predicted probabilities and the true labels.
@@
$model$ represents the parameters of the model. $Loss(model,\left(x^{i},y^{j}\right))$ represents the loss incurred by the model when predicting input $x^{i}$ with label $y^{j}$ from the poisoned dataset. $Loss(model,\left(x^{i^{\prime}},y^{j^{\prime}}\right))$ represents the loss incurred by the model when predicting input $x^{i^{\prime}}$ with label $y^{j^{\prime}}$ from the clean dataset. $n$ is the size of the poisoned dataset. $m$ is the size of the clean dataset.
@@ VII-B5 Pixel-space backdoor attack
Pixel-space backdoor attacks are a type of adversarial attack where an attacker modifies the pixel values of an image in a way that is imperceptible to the human eye but causes misclassification by the target model. A Pixel-space backdoor attack subtly modifies images in a way undetectable to the human eye but significantly alters the outcome of machine learning models. In the context of 6G, which heavily relies on AI-driven applications and ultra-reliable communications, such attacks could mislead autonomous systems or degrade the quality of service. The integrity and security of 6G-enabled applications, like autonomous vehicles or telemedicine, could be compromised, risking user safety and trust. Arshad et al.’s pixel-space backdoor attack [229] is a type of data poisoning attack where the attacker injects malicious samples into the training dataset to cause the model to learn a backdoor trigger. In this attack, the attacker flips certain bits of the pixel values of the training images, which can be done without significantly changing the image’s appearance. These modified images are then inserted into the training dataset, along with their original labels. During training, the model learns to associate the backdoor trigger with the target label, allowing the attacker to trigger the attack by adding the trigger to any input image. Arshad et al. evaluated their backdoor attack strategy on the MNIST digit dataset, which consists of 28x28 Grey-scale images of handwritten digits. They found that their attack was successful in 90% of the cases, meaning that the target model misclassified the poisoned images with the intended target label. Therefore, it’s important to note that the success of a backdoor attack depends on several factors, including the size and quality of the training dataset, the choice of the backdoor trigger, and the strength of the attacker’s evasion techniques. Therefore, while Arshad et al.’s attack achieved a high success rate on the MNIST dataset, it may not be as effective on other datasets or against more robust models with stronger defenses against adversarial attacks.
TABLE XV: Threat models of adversarial examples generation.
Attacks examples Attacks for the specific tasks White/blue/Grey box Adversarial Example Generation Methods
@@
Trained in a white-box or blue-box manner
@@
blue/Grey-box attack
@@
blue/Grey-box attack
@@
blue/Grey-box attack
@@
blue/Grey/White-box attack
@@
blue/Grey/White-box attack
@@
blue/Grey/White-box attack
@@ VII-C Adversarial examples
Adversarial examples refer to inputs specifically crafted to deceive machine learning models. Different adversarial attacks are presented, including blue-box attacks, white-box attacks, and Grey-box attacks. Akhtar and Mian [259] proposed a classification of adversarial examples on deep learning into two categories: 1) Attacks for Classification and 2) Attacks Beyond Classification/Recognition. The Attacks for Classification include miscellaneous attacks, adversarial transformation networks (ATNs), Houdini, Upset and ANGRI, Universal Adversarial Perturbations, Deepfool, Carlini and Wagner Attacks (C&W), One Pixel Attack, Jacobian-Based Saliency Map Attack (JSMA), Basic & Least-Likely-Class Iterative Methods, Fast Gradient Sign Method (FGSM), and Box-Constrained L-BFGS. Attacks Beyond Classification/Recognition include Attacks on Face Attributes, Semantic Segmentation & Object Detection, Deep Reinforcement Learning, Recurrent Neural Networks, and Attacks on Autoencoders and Generative Models. Table XV presents the threat models of adversarial examples generation. Figure 11 presents the adversarial examples in Edge Learning. Therefore, in our work, we propose a classification of adversarial examples into three categories: 1) White-box adversarial attacks, 2) blue-box adversarial attacks, and 3) Grey-Box adversarial attacks.
@@ VII-C1 White-box adversarial attacks
White-box attacks require complete knowledge of the targeted model, including its training method, architecture, parameter values, and, in most cases, its data training. A white-box adversarial attack on 6G networks exploits knowledge of the system’s architecture and algorithms to introduce crafted perturbations, undermining the accuracy and efficiency of data-driven applications. Such attacks can disrupt the ultra-reliable, low-latency communication fundamental to 6G, potentially compromising critical services and applications. To manipulate discrete text structure at its one-hot representation, Ebrahimi et al. [189] propose white-box adversarial examples, named HotFlip, which is applied for text classification. The HotFlip system uses the gradient concerning a one-hot input representation and can produce conflicting patterns with character substitutions ("flips"). The HotFlip also provides access to add and remove transactions by displaying them as character substitution sequences. The performance evaluation with AG’s news dataset and CharCNN-LSTM architecture shows that the adversary selects the flip transaction about 80% of the time and prioritizes suppression over insertion by two-to-one. Wang et al. [190] propose white-box attacks based on the integrated gradient for fooling deep neural networks. The proposed attacks are evaluated with the following three datasets: CIFAR-10, MNIST, and NSL-KDD. One of the most popular algorithms for generating white-box adversarial examples is the Fast Gradient Sign Method (FGSM) introduced by Goodfellow et al. in 2014 [253]. The FGSM generates adversarial examples by taking the gradient of the loss function concerning the input and adding a perturbation to the input in the gradient direction. The formula for generating adversarial examples using FGSM is as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{adv}=x+\epsilon\cdot sign(\nabla_{x}J(\theta,x,y))}$ (6)
@@
VII-C2 blue-box adversarial attacks
A blue-box adversarial attack targets 6G’s advanced machine learning algorithms, exploiting vulnerabilities to mislead or manipulate network functions. By subtly modifying input data, attackers can cause 6G systems to misinterpret or misclassify information, potentially compromising network efficiency, security, and reliability. Blue-box attacks provide a trained model with conflicting examples (during testing) produced without knowledge of the trained model. In particular cases, it is supposed that the adversary has restricted information about the model but does not know its parameters. Yang et al. [180] propose a blue-box method called TSadv, which attacks deep neural networks on time series. For solving the constrained optimization problem, the TSadv method uses a gradient-free attack strategy. The performance evaluation with the University of California Riverside (UCR) time series dataset shows good success rate, the average number of iterations, and mean squared error. Therefore, Biolková and Nguyen [184] propose a blue-box adversarial attack in speech recognition. The proposed attack uses a neural predictor that approximates the decision boundary’s length, resulting in a wrong audio signal transcription. The experiments on the LibriSpeech dataset show that the proposed attack can achieve a better success rate than BayesOpt [260] and SignOpt [261]. One of the most popular algorithms for generating blue-box adversarial examples is the Boundary Attack introduced by Brendel and Bethge in 2019 [262]. The Boundary Attack generates adversarial examples by iteratively exploring the decision boundary of the model. The formula for generating adversarial examples using Boundary Attack is as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{adv}=x+\alpha\cdot\frac{g}{\left|g\right|_{p}}}$ (7)
Where $x$ is the input, $\alpha$ is a small constant that controls the step size, $g$ is the gradient of the loss function concerning the input, and $\left|g\right|_{p}$ is the $l_{p}$ norm of the gradient.
@@ VII-C3 Grey-box adversarial attacks
A grey-box adversarial attack targets 6G networks by exploiting partial knowledge of the system’s architecture and parameters. Leveraging advanced machine learning techniques and the inherent complexities of 6G, attackers can introduce carefully crafted perturbations to disrupt network functionality. Grey-box adversarial attacks involve generating adversarial examples with limited knowledge of the targeted model. The attacker has access to some, but not all, of the model’s internal information. For example, the attacker may know the model’s architecture but not the weights or biases. The goal of a grey-box attack is to find an adversarial example that can deceive the targeted model into producing incorrect output, even though the attacker does not fully know the model’s internal workings. This is often done by leveraging the gradients of the model concerning the input, which can be computed even if the attacker does not know the exact values of the model’s parameters [263]. Grey-box attacks are particularly relevant in real-world scenarios, where attackers may not have access to the full details of a deployed model. However, they may still be able to mount effective attacks. Defenses against grey-box attacks typically involve techniques such as input sanitization [264] or model hardening, which aim to make it harder for attackers to find effective adversarial examples. One of the most popular algorithms for generating Grey-box adversarial examples is the Transferability Attack introduced by Papernot et al. in 2016 [265]. The Transferability Attack generates adversarial examples by exploiting the transferability property of adversarial examples, meaning that adversarial examples generated for one model can also fool another. The formula for generating adversarial examples using Transferability Attack is as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{adv}=x+\epsilon\cdot sign(\nabla_{x}J(\theta_{target},x,y))}$ (8)
This equation represents the process of generating an adversarial example for a given input image $x$, concerning a target class $\theta_{target}$ and a true label $y$, by adding a small perturbation $\epsilon$ to the original image $x$ in the direction of the sign of the gradient of the target class probability $J(\theta_{target},x,y)$ concerning the input image $x$.
In most cases, blue-box attacks are viewed as more difficult since the attacker has less information to operate with. Nevertheless, Grey box attacks can still be successful if the offensive party has sufficient knowledge of the pattern to produce successful adversarial examples.
@@ VII-D1 Synergetic attack
A synergetic attack consists of a combination of backdoor and adversarial examples against neural network classifiers. Synergetic attacks in 6G exploit the interdependence of heterogeneous technologies, combining vulnerabilities across multiple layers of the network. By targeting the intricate integration points in 6G’s collaborative systems, these attacks can disrupt seamless connectivity, compromising both service quality and security. Consequently, the complexity of 6G necessitates advanced security measures to counter such multifaceted threats. Liu et al. [179] proposed a synergetic attack, named AdvTrojan. The AdvTrojan is enabled based only on the model being infected through a backdoor during training, and the entries are thoughtfully disrupted. The AdvTrojan uses two stages that operate in parallel to progressively relocate the targeted input through the decision-making boundary to the adversary’s objective class. The model is injected with a Trojan backdoor during training in the initial stage. The Trojan backdoor is enabled during inference by incrementing the relevant target inputs using the pre-specified Trojan trigger. In the second stage, the targeted entry is completed by a prudent mixture of some adversarial disruption and the Trojan horse trigger. In the second stage, the targeted entry is completed by a prudent mixture of the Trojan horse trigger and some adversarial disruption. The adversary disruption enhances the Trojan horse trigger to exchange the entry label in the class the adversary targets. In practical terms, the Trojan horse trigger forwards the entry to an artificial position in the space of the entry closest to the boundary of the model’s decision. Afterward, the adversarial disruption makes the terminal push by shifting the forwarded example through the boundary of the decision, initiating the pre-established backdoor. The combined attacks can be defined as follows: